Hello!
Can somebody help please?
The members of a computer club were to be assessed equally to raise $400 for the purchase of software.
When four more persons joined the club, the per-member assessment was reduced by $5.00.
What was the new size of the club membership?
Thanks for your help and time
geomar

Answer by stanbon(75887)

Let "x" be the number of members before 4 were added.
Price per member is $400/x
After 4 members are added the number of members is "x+4"
Price per member is $400/(x+4)
EQUATION:
(per person price before) - (per person price after)=$5
400/x - 400/(x+4) = 5
Divide through by 5 to get:
80/x - 80/(x+4) = 1
Multiply thru by x(x+4) to get:
80(x+4) - 80x = x^2+4x
320=x^2+4x
x^2+4x-320=0
(x+20)(x-16)=0
x=16 (number of members before 4 were added)
Cheers,
Stan H.